# Communication-Efficient Soft Actor-Critic Policy Collaboration via Regulated Segment Mixture in Internet of Vehicles

### v2

@@ Abstract
Multi-Agent Reinforcement Learning (MARL) has emerged as a foundational approach for addressing diverse, intelligent control tasks, notably in autonomous driving within the Internet of Vehicles (IoV) domain. However, the widely assumed existence of a central node for centralized, federated learning-assisted MARL might be impractical in highly dynamic environments. This can lead to excessive communication overhead, potentially overwhelming the IoV system. To address these challenges, we design a novel communication-efficient and policy collaboration algorithm for MARL under the frameworks of Soft Actor-Critic (SAC) and Decentralized Federated Learning (DFL), named RSM-MASAC, within a fully distributed architecture. In particular, RSM-MASAC enhances multi-agent collaboration and prioritizes higher communication efficiency in dynamic IoV system by incorporating the concept of segmented aggregation in DFL and augmenting multiple model replicas from received neighboring policy segments, which are subsequently employed as reconstructed referential policies for mixing. Distinctively diverging from traditional RL approaches, with derived new bounds under Maximum Entropy Reinforcement Learning (MERL), RSM-MASAC adopts a theory-guided mixture metric to regulate the selection of contributive referential policies to guarantee the soft policy improvement during communication-assisted mixing phase. Finally, the extensive simulations in mixed-autonomy traffic control scenarios verify the effectiveness and superiority of our algorithm.
@@
Communication-efficient, Multi-agent reinforcement learning, Soft actor critic, Regulated segment mixture, Internet of vehicles.
@@ I Introduction
Internet of Vehicles (IoV) has evolved into a potent means to ubiquitously connect vehicles and enhance their self-driving capability. Various methods have been developed to control the speed of Connected Automated Vehicles (CAVs) in IoV to achieve better fleet management and accident avoidance, typically into three main types [2, 3, 4]: rule-based, optimization-based, and learning-based. Among these, learning-based Deep Reinforcement Learning (DRL) method has demonstrated superior driving performance [2, 5] and has become a widely discussed research area in IoV [6, 7, 8, 9, 10]. Specifically, CAVs can be modeled as DRL agents on top of a formulated Markov Decision Process (MDP). Correspondingly, these CAVs constitute a Multi-Agent Reinforcement Learning (MARL)-empowered system.
@@ I-A Problem Statement and Motivation
In general, most MARL works adopt a Centralised Training and Decentralised Execution (CTDE) architecture [11, 12, 13, 14]. The model training process in IoV scenarios can be conducted either in the cloud or on a MEC server’s Vehicular Digital Twins (VDT) [15, 16] or through on-device Vehicular Federated Learning (VFL) [17, 18, 19]. Specifically, benefiting from Federated Learning (FL) by its continuous learning capabilities and periodic model updates (i.e. gradients or parameters) exchange, VFL maintains sensitive information, such as location, driving habits, vehicle status, and control policy models locally, reducing latency and enabling instantaneous decision-making. However, the dynamic challenges, including high mobility, intermittent connectivity, localized and decentralized nature of vehicles, make the common Centralized Federated Learning (CFL)’s assumption of a super central training controller impractical and underlies potential threat to the stability and timeliness of overall learning performance [20]. Besides, generalization difficulties [21, 22], such as variations in physical-environment (e.g., road changes or weather conditions) and social-environments (interactions with Human-Driven Vehicles (HDVs)), necessitate continuous training or fine-tuning of decision-making models for specific environments or tasks with minimal on-line data requirements when deployment.
@@
In this way, there is growing research on peer-to-peer architecture’s Decentralized Federated Learning (DFL) [23] under VFL, focusing on the decentralized training/fine-tuning of Deep Neural Network (DNN) models among vehicles in real time [19, 24, 25]. This can be viewed as a distributed Stochastic Gradient Descent (SGD) optimization problem, in which the aggregation of exchanged DNN parameters acquired from periodic communication phase typically uses a simplistic parameter average mixture approach. In the context of Federated Reinforcement Learning (FRL) [26, 27, 28, 29], these exchanged and averaged parameters pertain to the RL agent’s policy network, which approximates the policy distribution through a DNN. And this can be implemented by Vehicle-to-Vehicle (V2V) communication as in most MARL works [30, 31, 32, 33]. However, the consequently frequent information exchange generates substantial communication overhead as the number of agents increases. Some DFL variants improve communication efficiency by reducing the number of communication rounds [34, 35, 36] or the communication workload per round [37, 24, 38, 39], but this can intensify the variability of local model updates, potentially leading to an inferior aggregated model post simple model parameter averaging [40]. This issue is more pronounced in online RL frameworks since RL agents interact more frequently with the environment than in offline supervised learning. The incremental arrival of data can magnify learning discrepancies among multiple agents and make vehicles less fault-tolerant. Interestingly, few studies have focused on the RL policies’ performance improvement after following this simple and straightforward aggregation of mixed policies using parameter averaging. However, not all communicated packets in MARL contribute effectively. During the communication-assisted mixing phase, MARL awaits a revolutionary policy parameter mixture method and corresponding metrics to regulate the aggregation of exchanged model updates, ensuring robust policy improvement for security and safety.
@@
On the other hand, there is a contradiction between limited and costly sample availability and the need for fast learning speeds in driving policy, that is, MARL algorithms must fully explore and adapt to dynamic environments with minimal online learning data requirements. This aligns with the principles of Maximum Entropy Reinforcement Learning (MERL), such as Soft $Q$-learning [41] and Soft Actor Critic (SAC) [42, 43], which transforms the traditional reward maximum into both expected return and the expected entropy of the policy. The incorporated entropy in optimization redefines value function and policy optimization objective from the ground up. This not only enhances sample efficiency, adaptability, and robustness in the dynamic and uncertain IoV environment, but also overturns the proof of the traditional performance improvement bound [44, 45, 46, 47]. Therefore, this work is dedicated to reanalyzing efficient communication within the MERL and DFL framework. We propose a practical policy mixture method underpinned by a rigorously derived mixture metric, enabling MERL to achieve smooth and reliable parameter updates not only during independent local learning phase but also throughout the communication-assisted mixing phase.
Figure 1: Illustration of MARL in autonomous driving.
@@ I-B Contribution
In this paper, we propose the Regulated Segment Mixture-based Multi-Agent SAC (RSM-MASAC) algorithm, tailored for training local driving models of CAVs in the IoV using MERL while addressing communication overhead challenges inherent in DFL. Our primary contributions include:
  • For the highly dynamic IoV setting, the proposed RSM-MASAC algorithm effectively combines communication-efficient DFL with MERL. Within RSM-MASAC, vehicles capably receive segments of policy networks’ parameters from neighbors within the communication range only and constitute referential policies for selective parameter mixture. Due to the well-designed policy parameter mixture method, RSM-MASAC ensures credible performance improvement during the communication-assisted mixing phase while maintaining sufficient exploration in the local learning phase.
@@
  • Through extensive simulations in the traffic speed control scenarios, our proposed RSM-MASAC algorithm could approach the converged performance of centralized FMARL [27] in a distributed manner, outperforming parameter average methods as in DFL [39, 24], thus confirming its effectiveness.
@@ I-C Related Works
In Multi-Agent Systems (MAS), Independent Reinforcement Learning (IRL) [53, 54, 55], which relies solely on agents’ local perceptions without any collaboration, has been extensively studied and often used as baselines due to its varied policy performance, unstable learning and uncertain convergence [27]. Examples of IRL methods include IQL [54], IAC [13], IA2C [56] and IPPO [57], etc.
@@
As an extension of IRL, distributed cooperative Reinforcement Learning (RL) enhances agents’ collective capabilities and efficiency by collaboratively seeking near-optimal solutions through limited information exchange with others. Notably, the exchanged contexts can be rather different and possibly include approximated value functions in [48, 49], rewards or even maximal $Q$-values on each state-action pair in [50]. Besides, given the clear evidence [58, 59] that experiences from homogeneous, independent learning agents in MAS can contribute to efficiently learning a commonly shared DNN model, the direct exchange of model updates during FL communication phase [26, 27, 60] is a viable way to indirectly integrate information and enhance cooperation among IRL. In addition to the commonly used centralized architecture that is less suitable for actual dynamic vehicular network, the peer-to-peer DFL paradigm, where clients exchange their local model updates only with their neighbors to achieve model consensus, emerges as an appealing alternative. This approach can be viewed as distributed SGD optimization, wherein the substantial communication expenditure cannot be overlooked. In that regard, some researchers resort to reduce communication frequency by aggregating more local updates before one round communication [34] or multi-round communication [35, 36]. Besides, either shrinking the number of parameters forwarded from local models or embarking on selective model synchronization is also tractable. For instance, Ref. [37] divides local gradients into several disjoint gradient partitions, with only a subset of these partitions being exchanged in any single communication round. Ref. [24] puts forward a randomized selection scheme for forwarding subsets of local model parameters to their one-hop neighbors. Furthermore, Ref. [38] suggests propagating top-k layers with higher normalized squared gradients, which may convey more information about the local data to neighbors. Meanwhile, Ref. [39] introduces a segmented gossip approach that involves synchronizing only model segments, thereby substantially splitting the communication expenditure.
@@
Moreover, while the above distributed SGD optimization methods and the works in actual VFL [17, 18, 19] assume ideal communication between cooperative agents, ensuring exact and perfect information sharing without compromising privacy concerns may not be feasible with suboptimal wireless transmission links. Apart from complex, customized, and costly cryptographic schemes [40, 61], most works [62, 51, 52] suggest local quantization of data before transmission, which prevents the exposure of raw data. Additionally, this quantization, along with compression or sparsification [63, 36, 64], can significantly reduce the message size, while the corresponding convergence analyses have been detailed in [51, 35, 52].
@@
Notably, when it comes to the method of mixing DNN parameters, these aforementioned DFL works generally adopt a simplistic averaging approach, which is familiar in parallel distributed SGD methods. However, when such an approach is directly applied into FRL, the crucial relationship between parameter gradient descent and policy improvement will be overshadowed [44]. Consequently, the corresponding mixture lacks proper, solid assessment means to prevent potential harm to policy performance [28]. In other words, directly using this kind of naive combination of communication efficient DFL and RL to enhance individual policy performance in IRL appears inefficient.
@@
Regarding this issue, we can draw inspiration from the conservative policy iteration algorithm [45], which leverages the concept of policy advantage as a crucial indicator to gauge the improvement of the cumulative rewards, and applies a mixture update rule directly for policy distributions in its pursuit of an approximately optimal policy. Moreover, the mixture metric utilized in the update rule is also investigated to avoid making aggressive updates towards risky directions. This is particularly relevant as excessively large policy updates often lead to a significant deterioration of policy performance [28]. TRPO [47] substitutes the mixture metric with Kullback-Leibler (KL) divergence, a measure that quantifies the disparity between the distributions of the current and updated policies, and enables agents to learn monotonically improving policies. Ref. [46] extends this work into cooperative MARL settings. However, directly mixing policy distributions is often an intractable endeavor. To implement this procedure in practical settings with parameterised policies, Ref. [44] further simplifies the KL divergence to the parameter space through Fisher Information Matrix (FIM), so as to improve the policy performance by directly mixing DNN parameters. It focuses on stable policy updates throughout the communication-assisted mixing phase, but its analysis is confined to the traditional policy iteration-based RL algorithm, whose primary objective revolves around maximizing the expected return.
@@
Transitioning to the realm of MERL, the integration of the entropy component marks a significant paradigm shift. MERL, exemplified by SAC, incorporates entropy maximization to encourage exploration. This dual objective of optimizing cumulative rewards and maintaining a high level of exploration fundamentally alters the criteria for policy improvement. As a result, traditional RL methodologies, particularly those based on PPO, are insufficient to address the new dynamics interplayed between reward maximization and exploratory behavior imposed by the entropy maximization in MERL. On the other hand, it’s also imperative to implement an appropriate and manageable mixture metric with monotonic policy improvement property to maximize the practicality of directly mixing policy parameters during the communication-assisted mixing phase. Such a metric must guarantee the efficacy of the resultant mixed policy in terms of policy improvement. Crucially, the assessment of the expected benefits of the mixed policy must be conducted before undertaking the actual policy mixing. That is, before any actual mixture of policy parameters, the potential of mixed policy to improve policy performance should be evaluated against the referential policy, whose parameters are received from neighbors in DFL. Only after confirming the anticipated benefits of the mixed policies should we proceed with the operation of mixing DNN parameters using the developed metric. Moreover, we have also summarized the key differences between our algorithm and the relevant literature in Table I. In conclusion, it is vital to reformulate the theoretical analysis within the combination of DFL and MERL framework, so as to offer robust theoretical underpinnings for practical applications in IoV.

### v3

@@ Abstract
Multi-Agent Reinforcement Learning (MARL) has emerged as a foundational approach for addressing diverse, intelligent control tasks in various scenarios like the Internet of Vehicles, Internet of Things, and Unmanned Aerial Vehicles. However, the widely assumed existence of a central node for centralized, federated learning-assisted MARL might be impractical in highly dynamic environments. This can lead to excessive communication overhead, potentially overwhelming the system. To address these challenges, we design a novel communication-efficient, fully distributed algorithm for collaborative MARL under the frameworks of Soft Actor-Critic (SAC) and Decentralized Federated Learning (DFL), named RSM-MASAC. In particular, RSM-MASAC enhances multi-agent collaboration and prioritizes higher communication efficiency in dynamic systems by incorporating the concept of segmented aggregation in DFL and augmenting multiple model replicas from received neighboring policy segments, which are subsequently employed as reconstructed referential policies for mixing. Distinctively diverging from traditional RL approaches, RSM-MASAC introduces new bounds under the framework of Maximum Entropy Reinforcement Learning (MERL). Correspondingly, it adopts a theory-guided mixture metric to regulate the selection of contributive referential policies, thus guaranteeing soft policy improvement during the communication-assisted mixing phase. Finally, the extensive simulations in mixed-autonomy traffic control scenarios verify the effectiveness and superiority of our algorithm.
@@
Communication-efficient, Multi-agent reinforcement learning, Soft actor-critic, Regulated segment mixture.
@@ I Introduction
Recently, Deep Reinforcement Learning (DRL) has gained significant traction in addressing complex, real-world applications, contingent on formulated Markov Decision Processes (MDPs) [2, 3]. Naturally, multiple collaborative DRL agents can form a scalable Multi-Agent Reinforcement Learning (MARL)-empowered system, efficiently coordinating sequential decision-making for complex scenarios [4, 5, 6, 7, 8]. Notable examples include fleet management and traffic control in the Internet of Vehicles (IoV), Unmanned Aerial Vehicles (UAV) and multi-robot systems, as well as distributed resource allocation in the Internet of Things (IoTs).
@@ I-A Problem Statement and Motivation
In general, most MARL works adopt a Centralised Training and Decentralised Execution (CTDE) architecture [9, 10, 11, 12]. Federated Reinforcement Learning (FRL) [13, 14, 15, 16] incorporates the continuous learning process in DRL with periodic model updates (i.e., exchanging gradients or parameters) from Federated Learning (FL), thus more competently coping with generalization difficulties during deployment [17, 18], such as variations in both physical and social environments. By keeping sensitive information localized, FRL enables instantaneous decision-making and promotes collaborative learning among independent agents. However, in most highly dynamic scenarios with high mobility, intermittent connectivity, and decentralized nature [19, 20], the common assumption of a super-centralized training controller in Centralized Federated Learning (CFL) becomes impractical, threatening both the stability and timeliness of overall learning performance [21]. Besides, generalization difficulties [22, 18], such as variations in external environments, necessitate continuous training or fine-tuning of decision-making models for specific environments or tasks with minimal online data requirements during deployment.
@@
In this way, there is growing research on peer-to-peer architecture’s Decentralized Federated Learning (DFL) [23], focusing on real-time decentralized training/fine-tuning of Deep Neural Network (DNN) models among DRL agents [24, 25, 26]. This can be viewed as a distributed Stochastic Gradient Descent (SGD) optimization problem, in which the aggregation of exchanged DNN parameters acquired from the periodic communication phase typically uses a simplistic parameter average mixture approach. In FRL, these exchanged and averaged parameters pertain to the DRL agent’s policy network, which approximates the policy distribution through a parameterized DNN. In addition, the parameter exchange can be implemented via Device-to-Device (D2D) or Vehicle-to-Vehicle (V2V) collaboration channels within communication range, common in many MARL works [17, 27, 7, 28]. However, the frequent information exchanges generate substantial communication overhead as the number of agents increases. Some DFL variants improve communication efficiency by reducing the number of communication rounds [29, 30, 31] or the communication workload per round [32, 25, 33, 34], but this can intensify the variability of local model updates, potentially leading to an inferior aggregated model post simple model parameter averaging [35]. This issue is more pronounced in online DRL frameworks since DRL agents interact more frequently with the environment compared to offline supervised learning. Incremental data can magnify learning discrepancies among agents and reduce fault tolerance. Interestingly, few studies have focused on the policy performance improvement issue of DRL with its combination with FL. Despite the simplicity and straightforward aggregation of mixed policies in parameter averaging, not all communicated packets in MARL contribute effectively. During the communication-assisted mixing phase, MARL awaits a revolutionary policy parameter mixture method and corresponding metrics to regulate the aggregation of exchanged model updates, ensuring robust policy improvement for security and safety.
@@
On the other hand, there is a contradiction between limited and costly sample availability and the need for fast learning speeds, that is, MARL algorithms shall fully explore and adapt to dynamic environments while minimizing the demand for extensive and costly online learning data. This aligns with the principles of Maximum Entropy Reinforcement Learning (MERL), such as Soft $Q$-learning [36] and Soft Actor-Critic (SAC) [37, 38], which transforms the traditional reward maximum into both expected return and the expected entropy of the policy. The incorporated entropy in optimization redefines value function and policy optimization objective from the ground up. Albeit its enhancement to sample efficiency, adaptability, and robustness in dynamic and uncertain environments, it overturns the proof of the traditional performance improvement bound [39, 40, 41, 42]. Among MERL algorithms, SAC [37, 38] is particularly well-suited for highly dynamic and uncertain environments due to multi-folded reasons. First, it leads to higher sample efficiency through off-policy learning. Second, it enables automatic entropy adjustment via optimization of the temperature parameter. Finally, by effectively combining value function learning with policy optimization, it significantly enhances stabilization and adaptability to continuous, complex, high-dimensional action space.
Therefore, this work is dedicated to reanalyzing efficient communication within the MERL and DFL framework. Prominently, a practical policy mixture method, which is underpinned by a rigorously derived mixture metric, is devised to enable SAC to achieve smooth and reliable parameter updates, not only during independent local learning phase but also throughout the communication-assisted mixing phase.
Figure 1: Illustration of Single-agent IRL and Multi-agent Distributed Cooperative RL.
@@ I-B Contribution
In this paper, we propose the Regulated Segment Mixture-based Multi-Agent SAC (RSM-MASAC) algorithm, tailored for training DRL agents under highly dynamic scenarios while addressing communication overhead challenges inherent in DFL. Our primary contributions include:
  • For the highly dynamic setting, the proposed RSM-MASAC algorithm effectively combines communication-efficient DFL with MERL. The algorithm enables agents to receive segments of policy networks’ parameters from neighbors within their communication range. These segments are used to constitute referential policies for strategically designed selective parameter mixture, ensuring credible performance improvement during the communication-assisted mixing phase while maintaining sufficient exploration in the local learning phase.
@@
  • Through extensive simulations in the traffic speed control task, a typical MARL-based IoV scenario, our proposed RSM-MASAC algorithm could approach the converged performance of centralized FMARL [14] in a distributed manner, outperforming parameter average methods as in DFL [34, 25], thus confirming its effectiveness.
@@ I-C Related Works
In Multi-Agent Systems (MAS), Independent Reinforcement Learning (IRL) [48, 49, 50], which relies solely on agents’ local perceptions without any collaboration, has been extensively studied and often used as a baseline due to its varied policy performance, unstable learning and uncertain convergence [14]. Examples of IRL methods include IQL [49], IAC [11], IA2C [51] and IPPO [52], etc.
@@
As an extension of IRL, distributed cooperative Reinforcement Learning (RL) enhances agents’ collective capabilities and efficiency by collaboratively seeking near-optimal solutions through limited information exchange with others, as illustrated in Fig. 1. Notably, the exchanged contexts can be rather different and possibly include approximated value functions in [43, 44], rewards or even maximal $Q$-values on each state-action pair in [45]. Besides, given the clear evidence [53, 54] that experiences from homogeneous, independent learning agents in MAS can contribute to efficiently learning a commonly shared DNN model, the direct exchange of model updates during FL communication phase [13, 14, 55] is a viable way to indirectly integrate information and enhance cooperation among IRL. In addition to the commonly used centralized architecture that is less suitable for actual dynamic environments, the peer-to-peer DFL paradigm, where clients exchange their local model updates only with their neighbors to achieve model consensus, emerges as an appealing alternative. This approach can be viewed as distributed SGD optimization, where the significant communication expenditure cannot be overlooked. In that regard, some researchers have developed strategies to reduce communication frequency by aggregating more local updates before one round communication [29] or multi-round communication [30, 31]. Besides, reducing the number of parameters transmitted from local models or implementing selective model synchronization is also tractable. For instance, Ref. [32] divides local gradients into several disjoint partitions, with only a subset being exchanged in any given communication round. Ref. [25] puts forward a randomized selection scheme for forwarding subsets of local model parameters to their one-hop neighbors. Furthermore, Ref. [33] suggests propagating top-k layers with higher normalized squared gradients, which may convey more information about the local data to neighbors. Meanwhile, Ref. [34] introduces a segmented gossip approach that involves synchronizing only model segments, thereby substantially splitting the communication expenditure.
@@
Moreover, while the above distributed SGD optimization methods and most works in actual IoV, IoT and UAV applications [19, 20, 24, 56, 4] assume ideal communication between cooperative agents, achieving exact and perfect information sharing without compromising privacy concerns may not be feasible with suboptimal wireless transmission links. Beyond complex, customized, and costly cryptographic schemes [35, 57], most research [58, 46, 47] suggests local quantization of data before transmission, effectively shielding raw data from exposure. Additionally, quantization, along with compression or sparsification [59, 31, 60], can significantly reduce the message size, with corresponding convergence analyses detailed in [46, 30, 47].
@@
Notably, when it comes to the method of mixing DNN parameters, these aforementioned DFL works generally adopt a simplistic averaging approach, which is familiar in parallel distributed SGD methods. However, when such an approach is directly applied to FRL, the crucial relationship between parameter gradient descent and policy improvement will be overshadowed [39]. Consequently, the corresponding mixture lacks proper, solid assessment means to prevent potential harm to policy performance [15]. In other words, directly using this kind of naive combination of communication efficient DFL and RL to enhance individual policy performance in IRL appears inefficient.
@@
Regarding this issue, we can draw inspiration from the conservative policy iteration algorithm [40], which leverages the concept of policy advantage as a crucial indicator to gauge the cumulative reward improvement and applies a direct mixture update rule for policy distributions in pursuit of an approximately optimal policy. Moreover, the mixture metric utilized in the update rule is also investigated to prevent overly aggressive updates towards risky directions, as excessively large policy updates often lead to significant policy performance deterioration [15]. TRPO [42] substitutes the mixture metric with Kullback-Leibler (KL) divergence, a measure that quantifies the disparity between current and updated policy distributions, facilitating the learning of monotonically improving policies. Ref. [41] extends this work into cooperative MARL settings. However, directly mixing policy distributions is often an intractable endeavor. To implement this procedure in practical settings with parameterized policies in DRL, Ref. [39] further simplifies the KL divergence to the parameter space through Fisher Information Matrix (FIM), so as to improve the policy performance by directly mixing DNN parameters. It focuses on stable policy updates throughout the communication-assisted mixing phase, but its analysis is confined to the traditional policy iteration-based RL algorithms, which aim to maximize the expected return only.
@@
Transitioning to the realm of MERL, the integration of the entropy maximization marks a significant paradigm shift, fundamentally redefining the criteria for policy improvement due to its dual objective of optimizing cumulative rewards and maintaining a high level of exploration. However, traditional RL methodologies, particularly those based on PPO, are insufficient to address the new dynamics interplayed between reward maximization and exploratory behavior imposed by the entropy maximization in MERL, as demonstrated in many works [61, 62, 37, 38]. On the other hand, it’s also imperative to implement an appropriate and manageable mixture metric with monotonic policy improvement property to maximize the practicality of directly mixing policy parameters during the communication-assisted mixing phase. Such a metric must guarantee the efficacy of the resultant mixed policy in terms of policy improvement. Crucially, the expected benefits of the mixed policy must be assessed before proceeding with actual policy mixing. Specifically, the mixed policy improvement should be evaluated against the referential policy, whose parameters are received from neighbors in DFL. Only after the anticipated benefits are confirmed should the operation to evaluate the established metric for mixing DNN parameters commence. Moreover, we have also summarized the key differences between our algorithm and relevant literature in Table I. In conclusion, it is vital to reformulate the theoretical analysis within the combination of DFL and MERL framework. More particularly, given its superiority among MERL algorithms, SAC [37, 38] is chosen as a showcase offering robust theoretical underpinnings for practical IoV, IoT, and UAV applications.